OpenStudy (anonymous):

multiply x^3/4 x x^1/5 x^4/9 x^3/20 x^11/20 x^19/20

OpenStudy (anonymous):

everything is multiplying so you add all those exponents

OpenStudy (anonymous):

When multiplying, exponents add. Thus \[x^{\frac 34} \times x^{\frac 15} = x^{\frac 34 + \frac15}\]\[= x^{\frac{3\times5+1\times4}{20}}\]\[=x^{\frac{19}{20}}\]